OpenStudy (anonymous):

enter equation of the following using line in slope intercept form (4,-2) (-2,4)

hartnn (hartnn):

first find slope, do u know how to find slope when 2 points are given ?

OpenStudy (anonymous):

no

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

hero (hero):

Alternatively, each point into y = mx + b: -2 = 4m + b 4 = -2m + b Then solve for m and b

OpenStudy (anonymous):

so how do i solve for m and b

hero (hero):

-2 = 4m + b 4 = -2m + b 1. Isolate b in each equation: -2 - 4m = b 4 + 2m = b 2. Set b = b: -2 - 4m = 4 + 2m 3. Place like terms on the same side: -2 - 4 = 4m + 2m 4. Combine like terms: -6 = 6m 5. Isolate m: -6/6 = m -1 = m 6. To solve for b, insert the value of m into one of the equations given in step 1: 4 + 2(-1) = b 4 -2 = b 2 = b 7. Input the values of m and b into the equation y = mx + b: y = -x + 2

OpenStudy (anonymous):

thank you